# Rethinking the effects of data contamination in Code Intelligence

### v2

@@ II-D2 Dataset preparation for LLMs:
To construct contaminated datasets and their uncontaminated counterparts for LLMs, we curate their released pre-training corpora, i.e., Google BigQuery for LLaMA and The Stack (v1.2) for StarCoder. The GitHub dataset available on Google BigQuery constitutes 4.5% of the pre-training corpus of LLaMA, accounting for 328 GB. As the main training base of StarCoder, Stack (v1.2) consists of 6.4 TB of source code in 384 PLs. We construct our experimental data from these two corpora. In Section II-E, we elaborate on the detailed procedures to construct contaminated datasets and their uncontaminated counterparts according to the specific requirements of each contamination setting.
Model | Research Question | Language | Experimental Group | Control Group | Perturbance | ||||
Avg | Std | Med | Avg | Std | Med | count (ratio) | |||
LLaMA | RQ1 | Java (JavaC#) | 62.47 | 27.09 | 55.50 | 62.84 | 26.88 | 57.00 | 9.73 (15.33%) |
C# (JavaC#) | 58.86 | 28.18 | 51.00 | 59.63 | 28.23 | 51.50 | 9.27 (15.49%) | ||
NL (JavaNL) | 30.99 | 7.56 | 30.00 | 30.36 | 7.31 | 30.00 | 15.54 (48.58%) | ||
RQ2 | NL (NLJava) | 90.49 | 22.59 | 86.00 | 99.22 | 29.07 | 95.00 | 27.58 (30.15%) | |
RQ3 | Java (JavaC#) | 60.09 | 61.00 | 38.50 | 59.94 | 60.95 | 37.50 | 12.41 (20.31%) | |
C# (JavaC#) | 61.13 | 68.37 | 40.50 | 61.13 | 68.50 | 39.00 | 13.20 (21.25%) | ||
RQ4 | Java (NLJava) | 47.75 | 46.03 | 32.00 | 47.99 | 45.27 | 31.50 | 9.20 (18.87%) | |
NL (NLJava) | 63.70 | 46.04 | 47.5 | 63.94 | 46.47 | 48.00 | 12.32 (19.04%) | ||
Starcoder | RQ1 | Java (JavaC#) | 52.30 | 21.20 | 48.00 | 52.82 | 21.01 | 49.00 | 8.21 (15.70%) |
C# (JavaC#) | 52.01 | 22.53 | 49.00 | 52.58 | 22.44 | 49.00 | 7.86 (15.11%) | ||
NL (JavaNL) | 28.02 | 6.18 | 27.50 | 28.26 | 6.23 | 28.00 | 14.26 (15.70%) | ||
RQ2 | NL (NLJava) | 86.77 | 23.10 | 81.00 | 96.66 | 25.78 | 90.50 | 26.66 (30.72%) | |
RQ3 | Java (JavaC#) | 38.62 | 28.29 | 27.50 | 38.97 | 28.57 | 27.00 | 5.48 (14.19%) | |
C# (JavaC#) | 38.84 | 28.86 | 28.00 | 39.15 | 28.96 | 28.50 | 5.82 (14.98%) | ||
RQ4 | Java (NLJava) | 28.71 | 13.35 | 26.00 | 29.51 | 13.52 | 26.50 | 6.33 (22.05%) | |
NL (NLJava) | 45.02 | 22.99 | 38.00 | 45.90 | 23.28 | 39.00 | 10.46 (23.23%) |
@@
Note: The Avg denotes the average number of tokens per sample, Std refers to the standard deviation of token counts, Med indicates the median number of tokens per sample, Language denotes the language of the input/output belongs to, and Perturbance count (ratio) represents the average number of perturbed tokens and their average proportion.
II-E Research Methodology
This section introduces the specific methodology that investigates the RQs proposed in Section II-A, especially the dataset construction for contaminated experimental groups and uncontaminated control groups.
II-E1 RQ1: Does Input-Only Contamination Lead to Performance Overestimation?
In this RQ, PLMs are experimented on code translation and generation tasks, while LLMs are evaluated on code translation and summarization tasks. Accordingly, we construct a series of exclusive datasets for different code tasks and PLMs/LLMs, dividing them into contaminated experimental groups and uncontaminated control groups.
@@ II-E1 RQ1: Does Input-Only Contamination Lead to Performance Overestimation?
Pre-trained Language Models: For the code translation task (JavaC#) of PLMs, we dictate each Java code as input while using each corresponding C# code as output. As for the uncontaminated control group, we directly use our prepared 160,000 Java samples for PLMs to pre-train from scratch, then fine-tune and infer on the CodeTrans dataset. Towards the contaminated experimental group, for each sample in CodeTrans’s testing set, we inject its Java portion into the PLMs’ pre-training corpus while leaving its C# portion intact. To eliminate the bias of pre-training size brought about by the contamination injection, an equal number of randomly selected samples is removed from the pre-training corpus. As for the code generation task (NLJava) of PLMs, the inputs are NLDs, while the outputs are Java code snippets. We follow the same procedure for the code translation task to configure the experimental setting.
@@
Large Language Models: For the code translation task (JavaC#) of LLMs, we first manually select 100 Java functions from both Google BigQuery’s GitHub and The Stack datasets as input portions of the contaminated experimental group. Afterwards, we manually migrate the above 100 Java functions into C#, forming code translation pairs, thereby ensuring the output portions are unseen to LLMs. To guarantee the migration quality, we follow Xue et al. [38]’s procedure and adapt their methodology into our manual migration, such as line-wise translation and keeping the naming conventions of C#. As for the code summarization task (JavaNL), we still adopt the above extracted Java functions as input portions, and manually craft their corresponding NLDs, thereby making the output portions (i.e., NLD) unseen to LLMs and forming the code summarization pairs of the contaminated experimental group. Similarly, we adhere to a series of standards when crafting NLDs. 1) NLDs must reflect the primary purpose of code snippets. 2) Limit NLDs to single sentences. 3) Focus on the high-level behaviors rather than low-level implementation details. Two authors with more than three years of Java and C# development experience are involved in the above manual procedures, one for implementation and another for double-checking.
Finally, to establish uncontaminated counterparts (i.e., control groups) in both code translation and summarization tasks, we perturb the above contaminated samples manually to ensure that both the input and output portions are unseen to LLMs. Specifically, the two authors participate in this procedure again. One is responsible for consistently perturbing both inputs and outputs, making sure they are still paired after perturbation. Besides, to guarantee that the task complexities are unchanged after perturbation for fair comparisons, only sub-token-level replacement is permitted (e.g., perturbing fieldName to fieldIdentifier), thereby maintaining that the code cyclomatic complexity [39] and code length are almost the same. After perturbation, another author double-checks the perturbation results to ensure the perturbation satisfies the consistency requirement mentioned above. To make readers learn more about our experimental datasets, we present Table I to illustrate their statistics, and a detailed illustration of each column is shown in the table note.
II-E2 RQ2: Does Output-Only Contamination Lead to Performance Overestimation?
In this RQ, both PLMs and LLMs are tested on code translation and generation tasks. The detailed experimental environment construction procedures are illustrated below.
Pre-trained Language Models: As for the code translation task (JavaC#) of PLMs, we follow all procedures in RQ1 to construct related experimental datasets. The only difference is that we use C# portion of CodeTrans’s testing set as contaminated data to inject into the pre-training corpora of both RoBERTa and GPT-2. As for the code generation task (NLJava), we still follow RQ1’s procedures except that we contaminate the Java portion (i.e., outputs) of CONCODE’s testing set.
Large Language Models: As for the experiments with LLMs, we take a different approach. Recall that we have prepared a series of input-output pairs for code translation and summarization tasks in RQ1. For the contaminated experimental group, their Java portions are extracted from Google BigQuery and Stack, which means they are contaminated. For simplicity, we directly swap the inputs and outputs for the contaminated experimental group and the uncontaminated control group, respectively, thereby constructing the output-only contamination setting. Thus, in Table I, statistics of C#Java of RQ2 are omitted. The above operations for the code translation task (C#Java) are enough owing to their line-wise equivalence. However, as for the code generation task (NLJava) derived from code summarization, to generate desired code snippets, previously crafted NLDs are too concise and lack implementation details. As such, we follow the criteria below to reconstruct NLDs for code generation. 1) Clearly state the input and output requirements. 2) Explicitly mention all external dependencies, such as libraries, classes, and interfaces. 3) Provide implementation-level details. Similarly, the previous two authors participate in this procedure again, one for implementation, another for double-checking. Specific statistics of experimental datasets are listed in Table I. As can be seen, the average token number of NL in RQ2 is larger than that in RQ1, owing to much more informative descriptions.
As for the uncontaminated control group, we follow the perturbation criteria mentioned in RQ1 to construct corresponding datasets, where their statistics concerning perturbation count and ratio are presented in Table I for reference.
II-E3 RQ3: Does Unpaired Contamination Lead to Performance Overestimation?
In this RQ, we still focus on code translation and generation tasks for PLM. Towards LLM, two dual code translation tasks are carried out.
Pre-trained Language Models: To conduct experiments with PLMs in this setting, we followed all the procedures from RQ1 to construct the datasets. The difference lies in the fact that we not only include the input portion of downstream tasks, but also add the corresponding output part to the pre-training dataset. Specifically, for the code translation task (JavaC#), we inject both the Java and C# portions of CodeTrans’s testing set into the pretraining corpus of PLMs separately, and remove an equal number of randomly selected samples from the pre-training corpus to eliminate the impact of the difference in training amount. Similarly, this process is also applied to the code generation task (NLJava).
Large Language Models: For conducting the unpaired contamination experiment on LLMs, we need to build an entirely new testing set with both input and output portions explicitly existing in the pre-training corpora of LLMs as independent samples. Focusing on the code translation task, we can further swap their input and output portions to form two dual code translation tasks (JavaC# and C#Java). Specifically, we use tree-sitter to filter Java and C# functions from Google BigQuery and Stack, respectively. Then, considering programs of different PLs with the same method names normally also carry similar or even identical functionalities, we transform their method names to lowercase for exact matching, thereby obtaining over one hundred million pairs of programs with identical names. To reduce the manual workload, we calculate the BLEU score between the Java and C# code snippets in the collected pairs and remove those with BLEU scores lower than 0.6, as we hypothesize that the program pairs that could serve as mutual translations should have relatively high similarities. Subsequently, we filter out program pairs exceeding 15 lines to reduce the generation difficulty for the LLMs, alleviating experimental errors caused by overly long tasks. After performing these two steps of filtering, we conduct manual screening on the remaining program pairs with the previous two authors again, one for program pair selection and another for double-checking. After 48 hours of manual screening, we obtain 100 unpaired contaminated samples from Google BigQuery and Stack, respectively, thereby constructing the contaminated experimental group for LLMs in this setting. For the uncontaminated control group, we followed the same perturbation process as in RQ1 for construction. Finally, via swapping the inputs and outputs for experimental and control groups, respectively, we form two dual code translation tasks (JavaC# and C#Java) satisfying unpaired contamination settings. Detailed statistics are shown in Table I.

### v3

@@ 2.4.2. Dataset preparation for LLMs:
To construct contaminated datasets and their uncontaminated counterparts for LLMs, we curate their released pre-training corpora, i.e., Google BigQuery for LLaMA and The Stack (v1.2) for StarCoder. The GitHub dataset available on Google BigQuery constitutes 4.5% of the pre-training corpus of LLaMA, accounting for 328 GB. As the main training base of StarCoder, Stack (v1.2) consists of 6.4 TB of source code in 384 PLs. We construct each code task’s experimental data from these two corpora, and the details can be found in Section 2.5.
| LLM | RQ | Task (In / Out) | Lang | Experimental Group | Control Group | Perturbance Count (Ratio) | |||||
| Avg | Std | Med | Avg | Std | Med | ||||||
| LLaMA | RQ1 | Code Translation (JaC#) | Ja | 62.47 | 27.09 | 55.5 | 62.84 | 26.88 | 57 | 9.73 (15.33%) | |
| C# | 58.86 | 28.18 | 51 | 59.63 | 28.23 | 51.5 | 9.27 (15.49%) | ||||
| Code Translation (PyJa) | Py | 80.61 | 32.43 | 73.5 | 81.64 | 31.96 | 75.5 | 11.53 (13.80%) | |||
| Ja | 123.1 | 58.71 | 115.5 | 123.09 | 59.48 | 108 | 17.96 (14.95%) | ||||
| Code Summarization (JaNL) | Ja | 62.47 | 27.09 | 55.5 | 62.84 | 26.88 | 57 | 9.73 (15.33%) | |||
| NL | 30.99 | 7.56 | 30 | 30.36 | 7.31 | 30 | 15.54 (48.58%) | ||||
| Code Summarization (PyNL) | Py | 80.61 | 32.43 | 73.5 | 81.64 | 31.96 | 75.5 | 11.53 (13.80%) | |||
| NL | 27.1 | 6.33 | 27 | 26.45 | 5.49 | 26.5 | 9.39 (33.85%) | ||||
| RQ2 | Code Translation (JaC#) | Ja | 93.83 | 45.31 | 84 | 90.26 | 42.93 | 81 | 20.27 (20.21%) | ||
| C# | 72.89 | 29.81 | 80 | 71.3 | 28.93 | 75.5 | 15.14 (19.50%) | ||||
| Code Translation (PyJa) | Py | 77.21 | 42.01 | 68.5 | 75.36 | 38.99 | 66 | 17.58 (21.23%) | |||
| Ja | 77.14 | 32.83 | 72.5 | 76.91 | 32.6 | 73 | 12.75 (17.50%) | ||||
| Code Generation (NLJa) | NL | 90.49 | 22.59 | 86 | 99.22 | 29.07 | 95 | 27.58 (30.15%) | |||
| Ja | 62.47 | 27.09 | 55.5 | 62.84 | 26.88 | 57 | 9.73 (15.33%) | ||||
| Code Generation (NLPy) | NL | 206.36 | 38.73 | 203 | 206.1 | 40.17 | 198 | 30.72 (14.89%) | |||
| Py | 80.61 | 32.43 | 73.5 | 82 | 31.99 | 76 | 11.54 (14.49%) | ||||
| RQ3 | Code Translation (JaC#) | Ja | 60.09 | 61 | 38.5 | 59.94 | 60.95 | 37.5 | 12.41 (20.31%) | ||
| C# | 61.13 | 68.37 | 40.5 | 61.13 | 68.5 | 39 | 13.20 (21.25%) | ||||
| Code Translation (PyJa) | Py | 29.34 | 16.59 | 25 | 29.23 | 15.74 | 25 | 5.5 (21.01%) | |||
| Ja | 42.78 | 20.81 | 38 | 42.45 | 20.08 | 38.5 | 9.33 (20.62%) | ||||
| RQ4 | Code Generation (NLJa) | NL | 47.75 | 46.03 | 32 | 47.99 | 45.27 | 31.5 | 9.20 (18.87%) | ||
| Ja | 63.7 | 46.04 | 47.5 | 63.94 | 46.47 | 48 | 12.32 (19.04%) | ||||
| Code Generation (NLPy) | NL | 50.94 | 41.53 | 37 | 52.62 | 43.63 | 37 | 6.42 (14,94%) | |||
| Py | 73.08 | 52.15 | 60.5 | 74.35 | 56.33 | 62 | 18.46 (25.26%) | ||||
| Code Summarization (JaNL) | Ja | 57.86 | 39.92 | 42.5 | 63.95 | 44.01 | 43.5 | 17.95 (31.02%) | |||
| NL | 18.46 | 6.54 | 18 | 19.8 | 6.9 | 19 | 4.51 (26.27%) | ||||
| Code Summarization (PyNL) | Py | 71.41 | 50.98 | 56 | 71.55 | 50.81 | 57 | 21.48 (30.08%) | |||
| NL | 16.6 | 5.64 | 16.5 | 17.79 | 5.94 | 17.5 | 4.11 (26.41%) | ||||
| StarCoder | RQ1 | Code Translation (JaC#) | Ja | 52.3 | 21.2 | 48 | 52.82 | 21.01 | 49 | 8.21 (15.70%) | |
| C# | 58.86 | 28.18 | 51 | 59.63 | 28.23 | 51.5 | 9.27 (15.49%) | ||||
| Code Translation (PyJa) | Py | 66.06 | 26.43 | 63 | 66.37 | 26.96 | 62.5 | 11.41 (17.86%) | |||
| Ja | 108.04 | 66.28 | 92.5 | 108.08 | 67.9 | 90 | 19.42 (17.69%) | ||||
| Code Summarization (JaNL) | Ja | 52.3 | 21.2 | 48 | 52.82 | 21.01 | 49 | 8.21 (15.70%) | |||
| NL | 28.02 | 6.18 | 27.5 | 28.26 | 6.23 | 28 | 14.26 (50.89%) | ||||
| Code Summarization (PyNL) | Py | 66.06 | 26.43 | 63 | 66.44 | 26.79 | 62.5 | 12.22 (19.26%) | |||
| NL | 25.23 | 4.71 | 25 | 24.51 | 4.37 | 24 | 10.24 (40.54%) | ||||
| RQ2 | Code Translation (JaC#) | Ja | 51.38 | 34.94 | 38.5 | 51.47 | 34.86 | 38.5 | 16.36 (31.84%) | ||
| C# | 32.92 | 18.58 | 29.5 | 32.91 | 20.61 | 30.5 | 10.52 (35.59%) | ||||
| Code Translation (PyJa) | Py | 64.03 | 43.4 | 53.5 | 64.32 | 45.27 | 55 | 15.43 (24.00%) | |||
| Ja | 54.92 | 33.16 | 47.5 | 54.55 | 34.6 | 47.5 | 12.95 (23.58%) | ||||
| Code Generation (NLJa) | NL | 86.77 | 23.1 | 81 | 96.66 | 25.78 | 90.5 | 26.66 (30.72%) | |||
| Ja | 52.3 | 21.2 | 48 | 52.82 | 21.01 | 49 | 8.21 (15.70%) | ||||
| Code Generation (NLPy) | NL | 216.5 | 44.25 | 206 | 222.01 | 45.45 | 209 | 22.67 (10.47%) | |||
| Py | 66.06 | 26.43 | 63 | 67.37 | 30.24 | 65 | 20.08 (30.40%) | ||||
| RQ3 | Code Translation (JaC#) | Ja | 38.62 | 28.29 | 27.5 | 38.97 | 28.57 | 27 | 5.48 (14.19%) | ||
| C# | 38.84 | 28.86 | 28 | 39.15 | 28.96 | 28.5 | 5.82 (14.98%) | ||||
| Code Translation (PyJa) | Py | 30.09 | 13.26 | 27 | 30.63 | 14.6 | 31 | 8.08 (29.05%) | |||
| Ja | 38.51 | 17.39 | 34 | 39.54 | 18.11 | 35.5 | 8.33 (21.63%) | ||||
| RQ4 | Code Generation (NLJa) | NL | 28.71 | 13.35 | 26 | 29.51 | 13.52 | 26.5 | 6.33 (22.05%) | ||
| Ja | 45.02 | 22.99 | 38 | 45.9 | 23.28 | 39 | 10.46 (23.23%) | ||||
| Code Generation (NLPy) | NL | 36.21 | 36.57 | 23 | 37.98 | 28.77 | 23.5 | 6.05 (19.08%) | |||
| Py | 77.21 | 47.64 | 69 | 76.85 | 52.73 | 69.5 | 25.68 (33.26%) | ||||
| Code Summarization (JaNL) | Ja | 59.94 | 36.2 | 53.5 | 62.15 | 36.92 | 53.5 | 13.03 (23.33%) | |||
| NL | 16.61 | 5.84 | 15 | 16.97 | 6.03 | 15 | 2.69 (17.31%) | ||||
| Code Summarization (PyNL) | Py | 72.27 | 55.85 | 59.5 | 71.32 | 65.02 | 58 | 20.06 (30.77%) | |||
| NL | 12.59 | 3.93 | 12 | 13.61 | 4.32 | 13 | 3.97 (33.49%) | ||||
@@
Note: The Avg denotes the average number of tokens per sample, Std refers to the standard deviation of token counts, Med indicates the median number of tokens per sample, Lang denotes the language of the input/output belongs to, and Perturbance count (ratio) represents the average number of perturbed tokens and their average proportion. Ja, Py, and NL denote Java, Python, and Natural Language, respectively. We use underlines to signify the contaminated portion of each RQ’s corresponding experimental group.
2.5. Research Methodology
This section introduces specific experimental procedures that investigate the RQs proposed in Section 2.1. In principle, both PLMs and LLMs should be evaluated on all three code tasks, i.e., code translation, generation, and summarization. PLMs indeed are the case, as their scales are feasible for pre-training from scratch, making it easier to precisely control the contamination settings. Nonetheless, since LLMs are too huge to be retrained from scratch, we cannot modify their pre-training corpora but have to proactively construct contaminated test sets according to their pre-training corpora. In this case, experimental settings for certain code tasks may not be satisfied and dropped. To make the whole narration clearer, we separately elaborate on the research methodology for PLMs and LLMs in each RQ. If certain code tasks are dropped in LLMs’ experiments, we will provide detailed explanations accordingly.
2.5.1. RQ1: Does Input-Only Contamination Lead to Performance Overestimation?
@@
(1) PLMs (Pre-train, fine-tune, then infer): For the Java side, code translation task (JavaC#) of PLMs dictating each Java code as input while using each corresponding C# code as output. When constructing the uncontaminated control group, we directly use our prepared 160,000 Java samples for PLMs to pre-train from scratch, then fine-tune and infer on the CodeTrans dataset. Towards the contaminated experimental group, for each sample in CodeTrans’s testing set, we inject its input portion (i.e., Java) into the PLMs’ pre-training corpus while leaving its output portion (i.e., C#) intact. To eliminate the bias of pre-training size brought about by the contamination injection, an equal number of randomly selected samples is removed from the pre-training corpus in advance in the experimental group. Afterwards, PLMs are pre-trained on the contaminated corpora, then fine-tune and infer on CodeTrans for performance comparison with the corresponding control group. As for the code generation task (NLJava), the inputs are NLs, while the outputs are Java code snippets, while the code summarization task (JavaNL) requires the Java code as inputs and NLs as outputs. We follow the same procedure as the code translation task to carry out the experiments, except that the fine-tuning and inference stages are supported by CONCODE and TL-CodeSum datasets, respectively.
In addition, for the Python-to-Java code translation, NL-to-Python code generation, and Python-to-NL code summarization tasks, we use the 250,000 Python samples for PLMs’ pre-training, while harnessing AVATAR, Gretel AI’s dataset, and code-docstring-corpus for their respective fine-tuning and inference. All other procedures are the same as the above Java side.
(2) LLMs (Pre-train, then infer): Starting from the Java side and its code translation task (JavaC#), we first manually select 100 Java functions from both Google BigQuery and The Stack datasets as input portions of the contaminated experimental group. Afterwards, we manually migrate the above Java functions into C#, forming code translation pairs, thereby ensuring the output portions are unseen to LLMs. To guarantee the migration quality, we follow Xue et al. (Xue et al., 2025)’s procedure and adapt their methodology into our manual migration, such as line-wise translation and keeping the naming conventions of C#. In practice, two authors with more than three years of Java, Python, and C# development experience are involved in this procedure above. One is responsible for implementation, while the other is for double-checking. As such, the experimental group of Java-to-C# code translation can be constructed. As for the uncontaminated control group, we perturb the above contaminated samples manually to weaken the model’s ability to directly memorize samples from pretraining. Besides, perturbations should be consistently applied to both samples’ inputs and outputs, making sure they are still paired after perturbation. Meanwhile, to guarantee that the task complexities (e.g., cyclomatic complexity (McCabe, 1976) and code length) are unchanged after perturbation for fair comparisons, only sub-token-level replacement in user-defined identifiers, e.g., function/variable names, is permitted (e.g., perturbing fieldName to fieldIdentifier), while leaving control flow structures, data types, and core logic intact. The two authors participate in this procedure again, one is responsible for the perturbation, while the other double-checks the perturbation results to ensure the perturbation satisfies the consistency requirement mentioned above. As such, we construct both the experimental and control groups for the Java-to-C# code translation task. Finally, we use LLMs directly conduct inference on their associated experimental and control groups for performance difference comparison.
@@
Regarding the code summarization task (JavaNL), we still adopt the above extracted 100 Java functions as contaminated input portions, and manually craft their corresponding NLs, thereby making the output portions (i.e., NL) unseen to LLMs and forming the code summarization pairs of the contaminated experimental group. Similarly, we adhere to a series of standards when crafting NLs. 1) NLs must reflect the primary purpose of code snippets. 2) Limit NLs to single sentences. 3) Focus on the high-level behaviors rather than low-level implementation details. The uncontaminated control group follows the same perturbation procedure as above for construction, thereby comparing performance with that of the experimental group. As for code tasks concerning Python, such as Python-to-Java code translation and Python-to-NL code summarization, we follow the same procedure to construct and compare both the experimental and control groups as above.
Code generation task (NLJava/Python) for LLMs’ experiments is ignored in this RQ, because it requires finding standalone NLs describing code functionalities without associated codes as contamination sources, in LLMs’ pre-training corpora, which are extremely scarce. To make readers learn more about the constructed datasets in both groups, we present Table 1 to illustrate their statistics, and a detailed illustration of each column is shown in the table note.
2.5.2. RQ2: Does Output-Only Contamination Lead to Performance Overestimation?
(1) PLMs (Pre-train, fine-tune, then infer): For the Java side, we follow the same experimental procedures as in RQ1, except that the contamination is injected on the output side rather than the input side. Specifically, for the Java-to-C# code translation task assessed on CodeTrans, we inject each of their testing samples’ output portion (i.e., C#) into the pre-training corpora of PLMs, while keeping the corresponding input portion (i.e., Java) intact. Similarly, for the NL-to-Java code generation and Java-to-NL code summarization tasks, we contaminate the pre-training corpus with the Java code and NL from their corresponding test sets, respectively, as they serve as the output side. The same strategy is also applied to Python-side evaluations, such as the Python-to-Java code translation, NL-to-Python code generation, and Python-to-NL code summarization tasks.
(2) LLMs (Pre-train, then infer): Still beginning with the Java side, the code translation task (JavaC#) follows a similar procedure to RQ1 for both experimental and control group construction and experimentation. The only difference is that we first randomly collect 100 C# code snippets from Google BigQuery and The Stack, then manually translate them into Java. As such, the input portion (i.e., Java) of this test set is unseen to LLMs, while the output portion (i.e., C#) is contaminated. For the code generation task (NLJava), we first randomly collect 100 Java code snippets from Google BigQuery and The Stack, then manually translate their functionalities into NL descriptions, making NL-to-Java pairs with the unseen inputs and contaminated outputs to LLMs. Other procedures, such as experimental and control groups construction, as well as experimentation, are the same as those in RQ1. The same procedure above is applied to code tasks on the Python side (PythonJava and NLPython). Besides, we do not conduct the code summarization task (Java/PythonNL) for LLMs in RQ2 owing to the same reason as RQ1’s code generation task.
2.5.3. RQ3: Does Unpaired Contamination Lead to Performance Overestimation?
(1) PLMs (Pre-train, fine-tune, then infer): To conduct experiments with PLMs under this setting, we follow all the procedures described in RQ1. The key difference is: when constructing contaminated experimental groups, we include not only the input portions of downstream tasks but also the corresponding output portions for contaminating PLMs’ pre-training corpora. For example, in the code translation task (JavaC#), we separate the Java and C# portions of each sample in the CodeTrans test set and randomly inject them into the pre-training corpora of PLMs, thereby constructing the contaminated experimental group. The same procedure is applied to code generation (NLJava) and summarization (JavaNL) tasks with their respective datasets on the Java side as well as to the equivalent Python-side experiments.